# Vehicle Rebalancing Under Adherence Uncertainty

### v1

@@ IV Case Studies
In this section, we discuss the experiments that we have performed to test the proposed driver state aware repositioning recommendation algorithm using large-scale fleet operations consisting of fleet size ranging from 2000 to 8000 taxis. We explain the operational details of the proposed driver preference and confidence-aware taxi repositioning recommendation model and discuss the components that will become the key ingredients of a realistic test-bed used for simulation case studies. We develop a simulated taxi network based on real data [23] from New York City. [^2] Specifically, we have used three months of data from [23] for training and testing the prediction models. The repositioning recommendation model (27) was triggered every $\mathcal{H}=60$ minutes in simulation. To generate the road network, we leveraged an approach similar to our prior work [8], in which city Manhattan was partitioned into regions denoted by the set $\mathcal{R}$.
@@
We first determine the centers of each region using the k-means clustering algorithm. Secondly, we leverage the road network to define a directed graph with the nodes representing the road junctions, and the edges representing the links between the road junctions. We then use Dijkstra’s shortest path algorithm to calculate the shortest distance route between the region centers. We denote $\mathcal{\delta}$ to be a matrix that stores the shortest path distance $\mathcal{\delta}_{ij}$ between region $i,j\in\mathcal{R}$. taxi drivers pick up passengers from one region and drop them off at the other region. After dropping off the customer, a driver has a choice to either keep waiting or reposition to a new region. Fig. 3 shows an idle-standing driver and the top 3 potential destinations which will maximize his likelihood of getting the ride in next $\mathcal{H}$ minutes.
@@ IV-A Taxi Driver’s Preference to Reposition
To evaluate the performance of the driver preference prediction problem, we use accuracy as the metric of evaluation which is defined in the context of top-$k$ destinations. We predict the top-$k$ destinations for a driver by selecting the regions with the highest $\mathcal{L}_{cj}$ values. If the driver’s actual choice of destination falls within these $k$ choices, we label the prediction as correct. The fraction of total number of correct predictions out of total number of predictions made for the driver is defined as the accuracy. As shown in right plot in Fig. 4, the accuracy with which the proposed model can pinpoint ($k=1$) the exact choice of driver destination is around 73%. Note that this accuracy varies from driver to driver due to a varied degree of randomness in each driver’s decision-making process and 73% is the average accuracy value for all the drivers. As the value of $k$ increases, the accuracy increases as well, e.g., we can predict that a driver will head to one of the two predicted destinations ($k=2$) with an accuracy of 83%, one of three predicted destinations ($k=3$) with an accuracy of 87% and one of four predicted destinations ($k=4$) with an accuracy of 90%. For the repositioning recommendation system, it is unnecessary to predict the exact destination of a driver, as this is neither possible with absolute certainty due to the inherent random choices of drivers, nor desirable, as we want drivers to remain more exploratory, i.e., they have a non-zero probability of moving to more than one region. This exploratory nature will ensure that the drivers can be steered to destinations in an optimized manner avoiding over supply. What is desired is a probability distribution over the regions, representing a driver’s choice of the repositioning destinations, which can be used to estimate the expected supply distribution (14).
@@
Figure 4: i)The figure on the left shows the feature importance plot highlighting that search distance plays the highest role in predicting taxi driver repositioning preference, ii) The figure on the right shows the accuracy improvement as the number of destination increases.
The dataset reported in [23] was used to extract the features for training the driver preference prediction model and Fig. 4 shows the features used for predicting the driver preference. As a first step, we need to identify the unique drivers in the dataset, which can be done by filtering out the drivers whose “medallion” matches with the “hack license”. A unique driver is the one for whom a trail of pick-ups and drop-offs which is coherent with time can be established. This means that every pickup is later than the previous drop-off. This information has been provided because in dataset [23], multiple drivers can have same “hack license” at the same time. Since, dataset [23] contains only the pick-up and drop-off data, a driver’s repositioning choice is defined as the next region of pickup after a drop-off. Each driver’s preference is learned by training a separate logistic regression model. The drivers whose choice can be predicted with the highest accuracy are selected. These drivers are chosen to be representative drivers, and a group of drivers are assumed to have the same preference as that of the representative driver in the group, e.g., if there are 2000 drivers in the simulation, then groups of 20 drivers may have same preference distribution as their representative driver. This assumption is taken because of practical limitation of data availability, e.g., in dataset [23] there are only 925 unique drivers and not all of them have a predictable preference. Each driver preference model is trained using the features shown in the left plot in Fig. 4. It was found that the top most feature to predict if a driver will move to the destination or not was the distance between the two regions. It is obvious that a driver will not go very far away in search of the passengers. Second influential feature is the median trip distance at the destination. This is because longer trip distance translates to higher profits. Note that the labels or targets in the logistic regression model are {1,0}, where 1 implies that the driver will go the destination. Since, raw data only contains the data points corresponding to the selected destinations (label=1), we augment the training data with label=0 for all the other destinations. The corresponding features such as search distance, median trip distance etc, can be easily extracted from the raw data. So, for each label=1 data point, we have 65 data points each with the label=0, since there are 66 regions in total. Once the logistic regression model is trained, we can evaluate its performance using accuracy evaluation, which measures how often the model correctly predicts {0,1}. However, this is not a good metric due to the significant imbalance between the label=0 and label=1 data points. The abundance of label=0 data points makes it easier to reject a potential destination, leading to misleadingly high accuracy values—around 99% for almost all drivers. So, we choose a modified accuracy as a metric of evaluation as detailed previously and shown in the right figure in Fig. 4 which plots the prediction accuracy vs number of predicted destinations.
@@
Figure 5: i) Plot on top left shows the distribution of the distance that a driver is willing to travel in search of passengers, ii) Plot on top right shows the distribution of the number of pickups that leads to the driver stay in the same region, iii) Plot on the bottom left shows the distribution of the trip distance at the destination iv) Plot on the bottom right shows the distribution of the hours in which the driver chooses to reposition or stay in the same region
@@
IV-B Taxi Driver’s Confidence on Recommender System
As the driver keeps interacting with the recommender system, the driver’s belief is updated using Eq. 8 in each time step. We create three scenarios based on the values of the weights $\epsilon_{0},\epsilon_{1}$:
  1. 1.
    Neutral $\epsilon_{0}=\epsilon_{1}$ : This means that the driver’s confidence will increase and decrease at the same rate due to successful as well as failed recommendations.
  2. 2.
    Pessimistic $\epsilon_{0}<\epsilon_{1}$ : This means that the driver’s confidence will decrease at a higher rate due to failed recommendations as compared to the rate of increasing in the confidence level due to successful recommendations.
  3. 3.
    Optimistic $\epsilon_{0}>\epsilon_{1}$ : This means that the driver’s confidence will increase at a higher rate due to successful recommendations as compared to the rate of decrease in the confidence level due to failed recommendations.
Fig. 6 shows how the confidence of a driver evolves over time when interacting with the recommender system. The confidence of the drivers in the optimistic case grows the fastest, followed by the neutral drivers, and slowest by the pessimistic drivers. Note that for the pessimistic drivers, the confidence level doesn’t converge to 1, rather it converges to 0.6. This means that there shall be some uncertainty at the steady-state in a driver’s response to recommendations. All the drivers initially start with a low confidence in the recommender system (0.2). In each time step, a driver is given a recommendation to reposition. If a driver is recommended to move to other destination, and the driver gets allocated within same time-step, then the recommendation is considered successful, else the recommendation is considered as failed.
Figure 6: This figure shows the evolution of the average confidence level of the drivers over time for three class of drivers.
IV-C Demand Prediction
@@
The performance evaluation of three demand forecasting models was conducted using key error metrics, as summarized in Table I. XGBoost demonstrated the best overall performance, achieving the lowest median absolute error of 3.91, compared to MOIRAI (5.03) and CHRONOS (8.49), indicating superior central tendency accuracy. At the 25th percentile, XGBoost also had the smallest error (0.97) followed by MOIRAI (1.33) and CHRONOS (1.98), showing better performance for relatively easier predictions. For the 75th percentile, XGBoost maintained its lead with an error of 11.02, outperforming MOIRAI (14.05) and CHRONOS (26.43), reflecting robustness for challenging predictions. At the 95th percentile, XGBoost continued to perform best, achieving an error of 31.79 compared to MOIRAI (40.66) and CHRONOS (77.29), highlighting its effectiveness in handling extreme outlier scenarios. Overall, XGBoost emerged as the most reliable model, followed by MOIRAI with moderate performance, while CHRONOS had the highest errors, indicating weaker predictive capability.
@@
Median Absolute Error 3.91 5.03 8.49
25th Percentile Error 0.97 1.33 1.98
75th Percentile Error 11.02 14.05 26.43
95th Percentile Error 31.79 40.66 77.29
@@
A spatial error analysis, summarized in Table II, further evaluates the forecasting performance across different regions. The analysis reveals the highest median errors (top 10 regions) ranging from 10.75 to 16.07 across regions, with percentage errors remaining relatively low, between 6.38% and 10.33%.
@@
IV-D Travel Time Prediction
The figure on the left in Fig. 8 shows the probability distribution of the error in travel time prediction. The median absolute error in predicting the inter-station travel time is approximately 2 minutes and the 99th percentile of error is approximately 11 minutes. The error in the travel time prediction impacts the reachability of the idle-sanding taxis as defined by Eq. 20. Reachability is defined as the fraction of inter region trips that can be served within the planning horizon (in this case 60 minutes) [5]. Duration of the planning horizon plays a critical role in determining the reachability of the taxi drivers. The figure on the right in Fig. 8 shows the reachability (percentage) vs planning horizon (minutes) plot. The black curve shows that a planning horizon approximately 30 minutes is sufficient to ensure that a taxi is able to reposition between any two regions within the planning horizon. Further, the red curve shows that in addition to repositioning, a taxi driver can also finish a trip if the planning horizon is further increased to 60 minutes. To achieve this, we only consider the trips that are no longer than 30 minutes. Such trips account for 95% of all the trips in the raw data. This helps us defining the time step to be 60 minutes in duration, and the benefits are two fold, first is that the planning horizon duration is sufficiently large so that the demands can be forecasted with high accuracy in terms of error percentage(small planning horizon duration lead to sparse time-series which are difficult to forecast) [30], and second is that it ensures that each taxi can reposition and serve a customer within the planning horizon, thus it makes it possible to discretize the simulation into time-steps, in which the sequence of operations in Algorithm 1 can be executed.
@@ IV-D Travel Time Prediction
IV-E Evaluation Metrics
To evaluate the performance of the proposed algorithm we perform a simulation over 7 days. Since each time step is 60 minutes in duration, the total number of time steps in the simulation are 168. For each time step, we record the information related to the performance of the repositioning recommender system and the following metrics are evaluated at the end of the simulation:
IV-E1 Allocation
Allocation is defined as the fraction of drivers that got allocated in a time step. If a driver is allocated in the time $k\in\mathcal{K}$ the reward $y_{c}(k)=\sum_{j\in\mathcal{R}}y_{cj}(k)$ is 1, else if the driver remains idle, then the reward is 0 as defined in (7). Allocation is evaluated using the following expression:
$\frac{1}{|\mathcal{K}|}\frac{1}{|\mathcal{C}|}\sum_{k\in\mathcal{K}}\sum_{c\in\mathcal{C}}y_{c}(k)$ (30)
IV-E2 Driver Profit
Profit for each driver $c\in\mathcal{C}$ is defined as the difference between the earnings made by the driver in time step $k\in\mathcal{K}$ and the rebalancing cost and allocation cost. As defined earlier, $x_{cj}(k)$ is the recommendation decision variable, and the earning of the driver $c\in\mathcal{C}$ be $\mathcal{P}_{cj}(k)$. Also, $y_{cj}(k)$ is the allocation decision variable and $\mathcal{Q}_{cj}(k)$ is the cost of moving the vehicle. We use the average driver profit as a metric which is evaluated as:
$\frac{1}{|\mathcal{K}|}\frac{1}{|\mathcal{C}|}\sum_{k\in\mathcal{K}}\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}y_{cj}(k)\cdot(\mathcal{P}_{cj}(k)-\mathcal{Q}_{cj}(k))-x_{cj}(k)\cdot\mathcal{Q}_{cj}(k)$ (31)
IV-E3 Driver Confidence
Median confidence of the drivers at the end of the simulation is used as a measure to evaluate the impact of the performance of the repositioning recommender system on a driver’s confidence which is evaluated as follows:
$Median(\mu(\mathcal{|\mathcal{K}|}))$ (32)
IV-E4 Met Demand
Met demand is the fraction of the total number of requests that were served. The served demand is also same the number of allocated taxis and is evaluated as:
$\frac{\sum_{k\in\mathcal{K}}\sum_{i\in\mathcal{R}}\sum_{c\in\mathcal{C}_{i}(k)}\sum_{j\in\mathcal{R}}y_{cj}(k)}{\sum_{k\in\mathcal{K}}\sum_{i\in\mathcal{R}}d_{i}(k)}$ (33)
TABLE III: Model Performance Analysis: Metric Evaluation
Fleet Size Metric Model (Optimistic) Baseline (Optimistic) Model (Neutral) Baseline (Neutral) Model (Pessimistic) Baseline (Pessimistic)
2000 Allocation 0.786 0.783 0.784 0.784 0.772 0.804
Driver Profit 11.242 10.065 11.231 10.094 11.076 10.519
Met Demand 0.374 0.373 0.373 0.373 0.368 0.383
Confidence 1.000 1.000 1.000 1.000 0.900 0.920
4000 Allocation 0.722 0.693 0.723 0.692 0.709 0.661
Driver Profit 9.742 9.020 9.748 9.043 9.640 8.800
Met Demand 0.687 0.659 0.688 0.659 0.675 0.629
Confidence 1.000 1.000 1.000 1.000 0.940 0.470
6000 Allocation 0.641 0.603 0.641 0.600 0.634 0.518
Driver Profit 8.440 7.973 8.453 7.946 8.422 6.951
Met Demand 0.915 0.861 0.915 0.857 0.904 0.740
Confidence 1.000 1.000 1.000 1.000 0.810 0.120
8000 Allocation 0.509 0.496 0.513 0.492 0.507 0.399
Driver Profit 6.725 6.518 6.771 6.486 6.737 5.347
Met Demand 0.969 0.944 0.976 0.936 0.966 0.760
Confidence 1.000 1.000 1.000 1.000 0.360 0.030
Figure 9: This figure shows the performance comparison of the Linearized OTRR (29) with baseline model [7] on metrics defined in IV-E.
IV-F Baseline Model
For comparing the performance of the proposed rebalancing model, we use the model by [7] as a baseline. This agent-level model has proved to be efficient in balancing supply-demand when the supply-agents are assumed to be absolutely adherent. It is agnostic to taxi driver’s preferences as well as impact of recommendations on the driver’s adherence to recommender system. Additionally, this model uses the same set of constraints as (27), so the primary difference is the two model is the objective function. The baseline optimization model is as follows:
@@ IV-F Baseline Model
Baseline model Waller et al. 2018 [7]
$\max_{x}$ $\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}x_{cj}\cdot\mathbb{E}[\mathbf{d_{j}}]\cdot(1-\frac{\mathcal{T}_{cj}}{\mathcal{H}})$ (34)
s.t. $\sum_{j\in\mathcal{R}}x_{cj}\leq 1\hskip 2.84544pt\forall c\in\mathcal{C}$

### v2

@@ IV Case Studies
In this section, we discuss the experiments that we have performed to test the proposed driver state aware repositioning recommendation algorithm using large-scale fleet operations. We explain the operational details of the proposed driver preference and confidence-aware taxi repositioning recommendation model and discuss the components that will become the key ingredients of a realistic test-bed used for simulation case studies. We develop a simulated taxi network based on real data [28] from New York City for the period of January to July 2010.222Specifically, we have used six months of data from [28] for training and one month for testing the prediction models. To generate the road network, we leveraged an approach similar to our prior work [8], in which Manhattan city was partitioned into regions denoted by the set $\mathcal{R}$.
@@
We first determine the centers of each region using the k-means clustering algorithm. Secondly, we leverage the road network to define a directed graph with the nodes representing the road junctions, and the edges representing the links between the road junctions. We then use Dijkstra’s shortest path algorithm to calculate the shortest distance route between the region centers. We denote $\mathcal{\delta}$ to be a matrix that stores the shortest path distance $\mathcal{\mathcal{D}}_{ij}$ between region $i,j\in\mathcal{R}$. taxi drivers pick up passengers from one region and drop them off at the other region. After dropping off the customer, a driver has a choice to either keep waiting or reposition to a new region. Fig. 4 shows an idle-standing driver and the top 3 potential destinations which will maximize his likelihood of getting the ride in next $\mathcal{H}$ minutes. The repositioning recommendation model (27) was triggered every $\mathcal{H}=5$ minutes to ensure rapid repositioning recommendation and avoid the scenarios where the taxi driver has to wait for the recommendation.
@@ IV-A Taxi Driver’s Preference to Reposition
[…]
@@
Figure 5: i)The figure on the left shows the feature weight plot highlighting that search distance plays the highest role in predicting taxi driver repositioning preference, ii) The figure on the right shows the accuracy improvement as the number of destination increases.
@@
Figure 6: i) Plot on top left shows the distribution of the distance that a driver is willing to travel in search of passengers, ii) Plot on top right shows the distribution of the number of pickups (in 30 mins) that leads to the driver stay in the same region, iii) Plot on the bottom left shows the distribution of the trip distance at the destination iv) Plot on the bottom right shows the distribution of the hours in which the driver chooses to reposition or stay in the same region
The dataset  [28] was used to extract the features for training the driver preference prediction model and Fig. 5 shows the features used for predicting the driver preference. As a first step, we need to identify the unique drivers in the dataset, which can be done by filtering out the drivers whose “medallion” matches with the “hack license”. A unique driver is the one for whom a trail of pick-ups and drop-offs which is coherent with time can be established. This means that every pickup is later than the previous drop-off. This information has been provided because in dataset [28], multiple drivers can have same “hack license” at the same time. Since, dataset [28] contains only the pick-up and drop-off data, a driver’s repositioning choice is defined as the next region of pickup after a drop-off. Each driver’s preference is learned by training a separate logistic regression model. The drivers whose choice can be predicted with the highest accuracy are selected. These drivers are chosen to be representative drivers, and a group of drivers are assumed to have the same preference as that of the representative driver in the group, e.g., if there are 2000 drivers in the simulation, then groups of 20 drivers may have same preference distribution as their representative driver. This assumption is taken because of practical limitation of data availability, e.g., in dataset [28] there are only 925 unique drivers and not all of them have a predictable preference. Each driver preference model is trained using the features shown in the left plot in Fig. 5. It was found that the top most feature to predict if a driver will move to the destination or not was the distance between the two regions. It is obvious that a driver will not go very far away in search of the passengers. Second influential feature is the median trip distance at the destination. This is because longer trip distance translates to higher profits. Note that the labels or targets in the logistic regression model are {1,0}, where 1 implies that the driver will go the destination. Since, raw data only contains the data points corresponding to the selected destinations (label=1), we augment the training data with label=0 for all the other destinations. The corresponding features such as search distance, median trip distance etc, can be easily extracted from the raw data. So, for each label=1 data point, we have 65 data points each with the label=0, since there are 66 regions in total. Once the logistic regression model is trained, we can evaluate its performance using accuracy evaluation, which measures how often the model correctly predicts {0,1}. However, this is not a good metric due to the significant imbalance between the label=0 and label=1 data points. The abundance of label=0 data points makes it easier to reject a potential destination, leading to misleadingly high accuracy values—around 99% for almost all drivers. So, we choose a modified accuracy as a metric of evaluation as detailed previously and shown in the right figure in Fig. 5 which plots the prediction accuracy vs number of predicted destinations.
@@
IV-B Demand Prediction
@@
The performance evaluation of three demand forecasting models was conducted using key error metrics, as summarized in Table I. XGBoost demonstrated the best overall performance, achieving the lowest mean absolute error of 4.05, compared to MOIRAI (6.47) and CHRONOS (29.99), indicating superior average accuracy. It also had the lowest root mean squared error (6.26), suggesting better stability in error distribution compared to MOIRAI (13.06) and CHRONOS (106.96). At the 25th percentile, XGBoost again led with an error of 0.90, followed by MOIRAI (1.32) and CHRONOS (3.81), showing its effectiveness in easier prediction scenarios. For the 75th percentile, XGBoost achieved an error of 5.54, outperforming MOIRAI (8.58) and CHRONOS (32.00), demonstrating robustness under moderately difficult conditions. At the 95th percentile, which highlights performance in extreme cases, XGBoost maintained its lead with an error of 13.23, ahead of MOIRAI (20.74) and CHRONOS (77.00). Overall, XGBoost emerged as the most reliable and consistent model, followed by MOIRAI with moderate performance, while CHRONOS exhibited the highest errors, indicating limited predictive capability.
@@
Mean Absolute Error 4.05 6.47 29.99
Root Mean Squared Error 6.26 13.06 106.96
25th Percentile Error 0.90 1.32 3.81
75th Percentile Error 5.54 8.58 32.00
95th Percentile Error 13.23 20.74 77.00
@@
A spatial error analysis, summarized in Table II, further evaluates forecasting accuracy across different regions. The analysis highlights the top 10 regions with the highest median errors, which range from 4.97 to 6.12, with percentage errors between 9.32% and 12.17%, indicating a relatively stable performance across the most challenging spatial zones.
IV-C Travel Time Prediction
The figure on the left in Fig. 8 shows the probability distribution of the error in travel time prediction. The median absolute error in predicting the inter-station travel time is approximately 2 minutes and the 99th percentile of error is approximately 11 minutes. The error in the travel time prediction as defined by Eq. 20 impacts the reachability of the idle-sanding taxis. Reachability is defined as the fraction of inter region combinations that can be traversed within the planning horizon [5]. Duration of the planning horizon plays a critical role in determining the reachability of the taxi drivers. The figure on the right in Fig. 8 shows the reachability (percentage) vs planning horizon (minutes) plot. The black curve shows that a planning horizon approximately 30 minutes is sufficient to ensure that a taxi is able to reposition between any two regions of Manhattan within the planning horizon. Further, the red curve shows that in addition to repositioning, a taxi driver can also finish a trip if the planning horizon is further increased to 60 minutes.
@@ IV-C Travel Time Prediction
IV-D Analysis of Adherence Aware Vehicle Rebalancing Problem
This section introduces the Linearized AAVR model and benchmarks it against four baseline, state-of-the-art rebalancing strategies. These baselines are implemented to enable a fair and structured comparison across agent-level and aggregate-level rebalancing formulations. The AAVR model uniquely incorporates driver-specific adherence behavior, while the baselines assume full compliance. Specifically, we compare our model (Problem IV.1) against four other ones: B1 [7], B2 as a single-period variant of [6], B3 [3], and B4 as the single-step variant of [18]. Models B1 and B2 are agent-level, producing driver-specific recommmendations directly. In contrast, B3 and B4 operate at the aggregate level and require a post-processing step (Problem IV.6) to generate driver-level actions.
We define the binary variable $\mathbf{x_{cj}}\in\{0,1\}$ as the decision variable for the recommendation at the driver level, where $\mathbf{x_{cj}=1}$ implies that driver $c\in\mathcal{C}$ is recommended to go to station $j\in\mathcal{R}$. Further, $\mathbf{X_{ij}}\in\mathbb{N}$ is defined as the total number of taxis recommended to go from station $i\in\mathcal{R}$ to $j\in\mathcal{R}$ and $\mathbf{Y_{ij}}\in\mathbb{N}$ defines the number of passengers in region $i\in\mathcal{R},$ that are allocated to the drivers in region $j\in\mathcal{R}$. Additionally, let $\mathbf{S}_{i}\in\mathbb{N}$ denote the total number of taxis present in region $i\in\mathcal{R}$ after the execution of rebalancing decisions. Let $\mathbf{Z}_{i}\in\mathbb{N}$ represent the number of taxis that successfully received allocations in region $i$, and $\mathbf{T}_{i}\in\mathbb{N}$ denote the number of unallocated passengers remaining in region $i$ following the rebalancing process. Finally, $\mathcal{D}_{cj}$ denotes the shortest path distance while $\mathcal{T}_{cj}$ denotes the shortest path travel time between driver $c$ and region $j$. The proposed model and the baseline optimization models B1-B4 are defined as follows:
@@ IV-D Analysis of Adherence Aware Vehicle Rebalancing Problem
AAVR (Linearized):
$\max_{\mathbf{x},\mathbf{Z}}$ $\sum_{j\in\mathcal{R}}\mathbf{Z_{j}}-\beta\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\cdot\mathcal{T}_{cj}$ (28)
s.t. $\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}=1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\mathbf{x_{cj}}\cdot\big{(}\mathcal{T}_{cj}-\mathcal{H}\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$\mathbf{Z_{j}}\leq\sum_{c}\mu(c)\cdot\mathbf{x_{cj}}+(1-\mu(c))\cdot L_{cj}\hskip 2.84544pt\forall j\in\mathcal{R}$
$\mathbf{Z_{j}}\leq\nu_{j}\hskip 2.84544pt\forall j\in\mathcal{R}$
The above model is the linearized equivalent of Problem II.1. A new decision variable $\mathbf{Z_{j}}\in\mathbb{N}$ is introduced to replace $\min\left(\mathbf{\mathbb{E}[\mathbf{d_{j}}],\mathbb{E}[\mathbf{s_{j}}]}\right)$ with bounds defined by the newly added constraints, where $\mathbb{E}[\mathbf{s_{j}}]=\sum_{c}\mu(c)\cdot\mathbf{x_{cj}}+(1-\mu(c))\cdot L_{cj}$ and $\mathbb{E}[\mathbf{d_{j}}]=\mathbf{\nu_{j}}$.
Problem IV.2.
Baseline Model 1 (B1)
$\max_{\mathbf{x}}$ $\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\cdot\mathbf{\nu_{j}}\cdot\left(1-\frac{\mathcal{T}_{cj}}{\mathcal{H}}\right)$ (29)
s.t. $\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\leq 1\hskip 14.22636pt\forall c\in\mathcal{C}$